Solve the system using substitution.
2x + 5y = -6
4x + y = -12
Please explain how to solve this.
Thank you

Answer by AnlytcPhil(1116)

Solve the system using substitution.
2x + 5y = -6
4x + y = -12
Please explain how to solve this. Thank you
--------------------------------------------------------
Solution: x = -3, y = 0 or the ordered pair (-3, 0)
--------------------------------------------------------
Here's a system just like it solved that you can use as a model for
solving yours:
6x + 7y = -42
7x + y = -6
Pick an equation and a letter in it to solve for. If possible, to make things
easier, pick an equation that contains a letter which has an understood
coefficient 1 or -1, and solve that equation for that letter.
So in this problem we pick the second equation to solve for y, since it has
an understood coefficient of 1 in the second equation:
7x + y = -6
y = -6 - 7x
Next enclose the expression for y, namely -4-7x, in parentheses, and
replace y by it in the equation that you haven't used yet.
6x + 7y = -42
6x + 7(-6-7x) = -42
6x - 42 - 49x = -42
-43x - 42 = -42
-43x = -42 + 42
-43x = 0
x = 0
That's the value for x. Now we must find the value for y. Substitute 0 for
y in the equation solved for y above:
y = -6 - 7x
y = -6 - 7(0)
y = -6 - 0
y = -6
So the answer is x=0, y=-6 or the ordered pair (0, -6)
Edwin